# Assessing AI vs Human-Authored Spear Phishing SMS Attacks: An Empirical Study Using the TRAPD Method

### v1

@@ 6.4 The TRAPD Methodology
There are several pros of the TRAPD methodology. First, it provides an ethical way for personalized deceptive messages to be evaluated, since it starts with targets giving consent to share their own personal information in ways that can be “weaponized,” while minimizing the effects of the deceptive messages. Second, the method provides data in a format that can statistically determine differences in effectiveness of different types of deceptive messages (e.g., AI vs Human, topic of message). Third, it allows targets to provide qualitative insights about their decision-making when ranking messages. It is difficult to collect such data from other methods (e.g., field experiments where targets receive a single message in an authentic environment). Furthermore, grounding the discussion in the actual messages themselves (e.g., “why did you place this message in the most likely to click on position?”), increases accuracy of their self-reported reasoning. Having them do this for multiple messages in one sitting also helps them recognize patterns. Finally, this method can be a positive learning experience for the targets who learn about their own thinking and what is most and least likely to deceive them.
@@
As with all methods, there are also drawbacks. These drawbacks primarily come from the fact that the targets provide self-reported data in a lab environment. For example, it is possible that participants say they would click on a link in a phishing message, when in fact they would not, or visa versa. Targets may have a social desirability bias, which would lead them to be less likely to admit falling for a deceptive message in the lab environment. Although it’s hard to know how strong this tendency may be, the majority of our participants (80%) said they would fall for at least one message. Furthermore, those who said they are confident in their ability to identify phishing messages said they would fall for fewer of them. It is also possible that individuals would act differently in more authentic contexts than what they expect when reporting in a lab. For example, a target may be less discerning when trying to quickly reply to a text message while riding a crowded bus than if they were at home on the couch. In our particular study, we used printouts of the text messages on the outline of a mobile phone to try and trigger similar thought patterns as they might have if looking on a real phone, but the lab environment and fact that it is paper and not digital was not possible to change. Additionally, this method focuses on the “content” of the messages, but cannot give input on some contextual factors, such as the source of the message. In the end, we have stronger confidence in the TRAPD method to identify differences between deceptive messages (and types of messages) than we do in the actual percentage of messages targets self-report as being willing to click on. Our observations of sessions with targets confirmed that they were fully engaged and took the sorting and threshold identification tasks very seriously.
@@
On a practical level, we learned several things about effectively implementing the TRAPD methodology that we hope will be useful to others in the future. We originally had planned on having targets sort 16 items, but after pilot testing, decided to lower the number to 12. Our observations suggest that it took participants about 10-15 minutes to sort 12. We would not recommend going above that number. As expected, not all targets who started the study came back for their interview. We ended up with a 61% return rate (25/41) after multiple requests for targets to return for the sorting interview. We propose that future studies plan for similar or possibly lower return rates. It is critical to make clear to targets that they are consenting to come back for an interview (if at all possible), not just to complete an initial survey. In this particular study, the final target count of 25 was low enough that we did not get statistically significant results in some areas that we likely would have with a larger sample. Because we already solicited the human generated messages that are targeted to the original targets, it is not practical to increase the sample without having to redo essentially the entire study. Thus the importance of getting a high return rate when using human-created messages. One other limitation of the current study, which can inform future studies, was the need to handle URLs consistently. We found that the GPT-4 generally created made-up links within messages, while many humans just included text such as “[URL]” or “link” instead of an actual link. We used a tinyurl.com link in such cases. Future studies may consider either using the same link within all messages, or enforcing the need for humans to create full URLs.
@@
The TRAPD methodology can be used for a variety of contexts where personalized deceptive messages are created and evaluated. Indeed, an early study evaluating the effectiveness of a spear phishing training approach versus a control group, used a less formally documented version of the TRAPD method [18]. We anticipate future studies using TRAPD to examine personalized disinformation, interactive AI-generated vishing attacks, and other forms of spear phishing. While using human-generated messages as part of TRAPD is clearly possible, there are advantages of using only automated messages. For example, imagine running a study that compares different LLMs or different prompts. The study could recruit targets, generate personalized deceptive content, and have targets rank order and describe their reasoning all within the same online survey. This is not possible with human-generated content, due to the needed delay to create the content, but would be possible with AI-only generated content. An online tool that supports TRAPD would likely help scale up the participant numbers.
@@ 7 Conclusion
Deceptive messages that are personalized to a particular individual, such as spear phishing messages, can be highly persuasive. We are entering a new age, when LLMs can be used to generate personalized deceptive messages. This study uses a novel methodology, we call TRAPD, to compare the efficacy of deceptive spear phishing SMS messages created by humans and AI. Although our statistical results are not definitive (due to a sample size of 28), we find a high likelihood (80%) that AI outperforms humans. We also find that messages related to jobs outperform those related to hobbies and items purchased online. Participants who were targeted with spear phishing messages described reasons they classified personalized smishing messages as particularly deceptive or easy to identify as fake. These were classified into the following key categories, in order of importance: characteristics of the URL, proper use of the technology communications medium of texting, use of inaccurate information, degree of personal relevance, application of the scarcity principle, messaging style, use of plausible rewards, and characteristics of the stated sender. We also report on the failure of targets to identify which messages are created by AI versus humans. Targets were not confident in their ability to identify AI-generated content. They were often inconsistent in their assessment of the features that they thought indicated a message was AI-generated. Features that influenced their AI vs human message decisions included the message having emojis, characteristics of the message URL, word choice, message style, length of messages, structure of messages, grammar/spelling, and if and how personalization was used. In summary, current LLMs can create highly deceptive spear phishing messages personalized to a target without targets having any idea that they are created by AI. This poses significant risks for cybersecurity breaches and societal resilience.

### v2

@@ 6.4 The TRAPD Methodology
There are several pros of the TRAPD methodology. First, it provides an ethical way for personalized deceptive messages to be evaluated, since it starts with targets giving consent to share their own personal information in ways that can be “weaponized," while minimizing the effects of the deceptive messages. Second, the method provides data in a format that can statistically determine differences in the effectiveness of different types of deceptive messages (e.g., AI vs Human, topic of message). Third, it allows targets to provide qualitative insights about their decision-making when ranking messages. It is difficult to collect such data from other methods (e.g., field experiments where targets receive a single message in an authentic environment). Furthermore, grounding the discussion in the actual messages themselves (e.g., “why did you place this message in the most likely to click on position?") supports accurate self-reported reasoning. Having them do this for multiple messages in one sitting also helps them recognize patterns. Finally, this method can be a positive learning experience for the targets as they evaluate their own thinking and what is most and least likely to deceive them.
Survey methods are often criticized for lower ecological validity compared to field experiments. However, they remain effective for understanding participants’ perceptions and controlled comparisons. Prior research suggests, particularly in phishing studies, they remain valid in capturing behavioral variables, like participants’ perceptions of phishing messages[43]. Research also indicates that contextual factors can significantly influence participants’ behavior toward phishing attacks[44]. We strengthen TRAPD’s ecological validity by incorporating personalized messages tailored to participants’ own contexts, thus aligning their evaluations more closely with real-world decision-making. Although field studies are generally more ecologically valid, they can raise ethical concerns as studies of this nature often involve deception and getting permission to create personalized phishing messages without tipping off users by gaining consent can be difficult. TRAPD addresses both these concerns by using an adequately valid method of collecting data in a controlled, ethical environment.
@@
As with all methods, there are also drawbacks. These drawbacks primarily come from the fact that the targets provide self-reported data in a lab environment. For example, it is possible that participants say they would click on a link in a message, when in fact they would not, or vice versa. Targets may have a social desirability bias, which would lead them to be less likely to admit falling for a deceptive message in the lab environment, if they recognize that the messages are deceptive in nature. Although it is hard to know how strong this tendency may be, the majority of our participants (80%) said they would click on the link in at least one message. Furthermore, those who said they were confident in their ability to identify phishing messages (in the questionnaire a month prior to the sorting/interview activity) said they would fall for fewer of them. It is also possible that individuals would act differently in more authentic contexts than what they expect when reporting in a lab. For example, a target may be less discerning when trying to quickly reply to a text message while riding a crowded bus than if they were at home on the couch. In our particular study, we used printouts of the text messages on the outline of a mobile phone to try and trigger similar thought patterns as they might have if looking on a real phone, but the lab environment and fact that it is paper and not digital was not possible to change. Additionally, this method focuses on the “content" of the messages, but cannot give input on some contextual factors, such as the source of the message. In the end, we have stronger confidence in the TRAPD method to identify differences between deceptive messages (and types of messages) than we do in the actual percentage of messages targets self-report as being willing to click on. Our observations of sessions with targets confirmed that they were fully engaged and took the sorting and threshold identification tasks very seriously.
@@
On a practical level, we learned several things about effectively implementing the TRAPD methodology that we hope will be useful to others in the future. We originally had planned on having targets sort 16 items, but after pilot testing, decided to lower the number to 12. Our observations suggest that it took participants about 10-15 minutes to sort 12. We would not recommend going above that number. As expected, not all targets who started the study came back for their interview. We ended up with a 61% return rate (25/41) after multiple requests for targets to return for the sorting interview. We propose that future studies plan for similar or possibly lower return rates. It is critical to make clear to targets that they are consenting to come back for an interview (if at all possible), not just to complete an initial survey. In this particular study, the final target count of 25 was low enough that we did not get statistically significant results in some areas that we likely would have with a larger sample. Because we already solicited the human generated messages that are targeted to the original targets, it is not practical to increase the sample without having to redo essentially the entire study. Thus the importance of getting a high return rate when using human-created messages. One other limitation of the current study, which can inform future studies, was the need to handle URLs consistently. We found that the GPT-4 generally created made-up links within messages, while many humans just included text such as “[URL]" or “link" instead of an actual link. We used a tinyurl.com link in such cases. Future studies may consider either using the same link within all messages or enforcing the need for humans to create full URLs.
@@
The TRAPD methodology can be used for a variety of contexts where personalized deceptive messages are created and evaluated. Indeed, an early study evaluating the effectiveness of a spear phishing training approach versus a control group used a less formally documented version of the TRAPD method [45]. We anticipate future studies using TRAPD to examine personalized disinformation, interactive AI-generated vishing attacks, and other forms of spear phishing. While using human-generated messages as part of TRAPD is clearly possible, there are advantages of using only automated messages. For example, imagine running a study that compares different LLMs or different prompts. The study could recruit targets, generate personalized deceptive content, and have targets rank order and describe their reasoning all within the same online survey. This is not possible with personalized human-generated content, due to the needed delay to create the personalized content, but would be possible with AI-only generated content. An online tool that supports TRAPD would likely help scale up participant numbers.
6.5 Implications of AI-enabled Spear Phishing
Currently, generative AI use is no longer a novelty among enthusiasts and researchers. As of 2024, a global survey by McKinsey & Company reported that 72% of organizations have adopted AI in at least one business function .[46] Despite the myriad of positive, legitimate uses that generative AI offers, there is increasing concern about the ethical challenges of its potential misuse. Particularly, there is an inherent risk that malicious actors can leverage these powerful models to be repurposed for spear phishing. With how effective LLMs can provide output through prompts that do not require much coding skill, this can potentially allow users who are not particularly technologically savvy to prompt AI to create spear phishing content. Accountability is also a big concern when dealing with AI abuse—should the developers also be accountable if their AI was used for nefarious uses? The language of existing laws regarding cybercrime are often aimed towards traditional phishing methods and may not be enough to address the unique challenges of AI.[47] A deeper exploration on the capabilities of AI and the safeguards required to prevent misuse will be a significant motivation for policymakers to promote discourse and implement standards.
Despite recent advancements in the field, there is still a significant gap in literature regarding AI-enabled spear phishing. Cybersecurity is an arms race between attackers and defenders, thus research to better understand these risks is essential. While many studies offer updated training methods to inform employees and end-users about phishing attacks [45], this study aligns with these efforts while also addressing the emerging threat of AI-generated spear phishing. While several studies have examined tools, frameworks, and AI or machine learning capabilities to detect phishing messages [48, 49, 50], little is known how publicly available LLMs can succeed in this endeavor. By evaluating how effectively publicly available LLMs can generate convincing spear phishing attacks, we can emphasize the need for countermeasures for potential malicious AI use. We are hopeful that this study will bridge the gap in the existing literature.
Understanding whether a phishing message was created by AI or a human has broader implications for cybersecurity policy. AI-generated messages are cheaper and faster to produce, but human-generated messages can introduce unique context and adaptability that AI might lack. Our findings suggest that distinguishing AI- from human-generated phishing messages is not merely an academic exercise but a practical necessity for developing tools that can detect and counteract evolving threats. Future studies should focus on refining this distinction and exploring how attackers might blend AI and human strategies to create even more convincing messages.
As the results from this study suggest, derivative studies may explore methods to enhance cybersecurity training using personalized spear phishing messages generated by AI. Tailoring training exercises to reflect realistic and specific phishing scenarios can improve user awareness and resilience. Future research could also investigate the psychological and behavioral factors that contribute to social engineering susceptibility. Additionally, collaborations between policymakers, AI developers, and cybersecurity experts are crucial to establish safeguards that mitigate AI misuse while preserving innovation. Research efforts could also focus on refining AI detection systems to better identify and flag AI-generated phishing attempts. Lastly, studies examining the long-term impact of these implications could provide actionable insights to shape global cybersecurity standards.
7 Limitations
This section describes the various limitations that exist in this pilot study:
7.1 Phishing Message Design
The phishing messages were crafted by students, which may not match the sophistication of the messages created by professional social engineers. Nevertheless, the messages were validated and reviewed by cybersecurity professors to ensure an acceptable quality of the human messages. Furthermore, for each prompt, three student-generated messages were evaluated, and the two most convincing were selected. The phishing messages deliberately lacked contextual cues, such as the senders’ email address or URL, to isolate and evaluate the effectiveness of the content itself, without the influence of other factors. This study would not factor in responses by participants when encountering suspicious senders or links. While it may affect the number of participants who would or would not click on a link, it is unlikely to affect the relative rankings of the messages based on their content.
7.2 Study Design
Our methodology used a self-reported approach to measure susceptibility, which may not accurately reflect actual behavior, as participants might downplay or overemphasize their susceptibility. This is not an uncommon methodology, as there are studies that use self reporting to assess convincingness of a message .[51, 52, 53] This method was also chosen to ensure the study remained ethical and avoided exposing participants to real spear phishing attacks or privacy violations. This method also allowed for direct comparison between messages which was useful to answer our research questions. Also, the relatively small sample size (25) limits the generalizability of the findings and statistical power of the study. However, as this is a pilot study, our primary purpose is to serve as a stepping stone for future research.
@@ 8 Conclusion
Deceptive messages that are personalized to a particular individual, such as spear phishing messages, can be highly persuasive. We are entering a new age, when LLMs can be used to generate personalized deceptive messages. This study used a novel methodology, we call TRAPD, to compare the efficacy of deceptive spear phishing SMS messages created by humans and AI. Although our statistical results are not definitive (due to a sample size of 25), we find a high likelihood (80%) that AI outperforms humans. We also find that messages related to jobs outperform those related to hobbies and items purchased online. This study highlights the capability of current-gen AI to replicate or even surpass humans in crafting convincing spear phishing messages.
Participants who were targeted with spear phishing messages described reasons they classified personalized smishing messages as particularly deceptive or easy to identify as fake. These were classified into the following key categories, in order of importance: characteristics of the URL, proper use of the technology communications medium of texting, use of inaccurate information, degree of personal relevance, application of the scarcity principle, messaging style, use of plausible rewards, and characteristics of the stated sender. These findings reinforce the importance of understanding the attributes of personalized deceptive messages, in order to develop countermeasures.
We also report on the failure of targets to identify which messages are created by AI versus humans. Targets were not confident in their ability to identify AI-generated content. They were often inconsistent in their assessment of the features that they thought indicated a message was AI-generated. Features that influenced their AI vs human message decisions included the message having emojis, characteristics of the message URL, word choice, message style, length of messages, structure of messages, grammar/spelling, and if and how personalization was used. This inability to reliably distinguish AI- from human-generated messages presents a critical challenge for cybersecurity. As AI-generated content becomes increasingly indistinguishable from human-authored material, individuals may rely on outdated heuristics, such as grammar errors, that are no longer applicable.
In summary, current LLMs can create highly deceptive spear phishing messages personalized to a target without targets having any idea that they are created by AI. Future research should focus on refining detection technologies and improving end-user training to address the evolving threat of AI-enabled spear phishing. By doing so, we can better anticipate and mitigate these emerging risks.